# Prospects for Probing Sub-GeV Leptophilic Dark Matter with the Future VLAST

### v1

@@ III.1 Benchmark models
In Ref. Abdughani et al. (2022), all representative interactions for muonphilic DM models were systematically classified. These models include DM and mediator (MED) particles, which can be scalar, fermion, or vector. To prevent DM decay, we introduced a $Z_{2}$ symmetry and constructed 16 renormalizable effective models with $Z_{2}$-even mediators, as summarized in Table II of Ref. Abdughani et al. (2022).
In this study, we first perform a preliminary scan of these 16 models and categorize them into three groups based on their annihilation characteristics: $s$-wave, $p$-wave, and $s$+$p$-wave. Moreover, we find that models exhibiting the same annihilation behavior share highly similar surviving parameter spaces after incorporating the experimental constraints listed in Table 1. Given these findings, we select three benchmark models as the focus of our analysis and presentation: scalar DM with a scalar mediator, Dirac DM with a scalar mediator, and Dirac DM with a vector mediator. The corresponding Lagrangians for these models are denoted as $\mathcal{L}_{SS}$, $\mathcal{L}_{FS}$, and $\mathcal{L}_{FV}$, respectively. The detailed discussions of these models is presented in the following section.
For DM particles promptly annihilating into detectable $\gamma$-rays with energy $\mathcal{O}(10\ \text{MeV})$, their mass must be in the MeV to GeV range. Additionally, studies Matsumoto et al. (2019); Chen et al. (2024, 2018) indicate that the constraints of the thermal relic density requires sub-GeV DM annihilation via a mediator with mass below GeV. For leptophilic mediators, such a light particle can only decay into muons or electrons. Therefore, to focus on leptophilic mediators, we individually study two annihilation channels: (i) ${\rm DM}+{\rm DM}\to e^{+}+e^{-}$ or ${\rm DM}+{\rm DM}\to 4e$, and (ii) ${\rm DM}+{\rm DM}\to\mu^{+}+\mu^{-}$ or ${\rm DM}+{\rm DM}\to 4\mu$. Based on these motivations, we list three representative interactions.
@@
    where $f$ can be either $e^{\pm}$ or $\mu^{\pm}$ in this work. The dimensional coupling between $S$ and $\phi$ is $M_{D}$, while $g_{f}$ denotes the coupling between the mediator and the fermions. This interaction suggests that DM annihilation proceeds via $s$-wave, with a cross-section around $\mathcal{O}(10^{-26})$ cm-3s-1 to produce the correct relic density $\Omega h^{2}\approx 0.1$. However, such $s$-wave annihilation with DM masses below $1\,\mathrm{GeV}$ may inject excessive energy, potentially distorting the temperature and polarization spectra of the CMB Matsumoto et al. (2019); Chen et al. (2024); Wang et al. (2025). Therefore, CMB observations impose stringent constraints on the coupling constants $g_{D}$ and $g_{f}$, particularly for sub-GeV DM $s$-wave annihilation.
@@ III.2 Experimental Constraints
Likelihood Constraints
Relic abundance Gaussian $\Omega_{\rm{DM}}^{\rm exp}h^{2}=0.1193\pm 0.0014$ Aghanim et al. (2020a); $\sigma_{\rm sys}=10\%\times\Omega_{\rm{DM}}^{\rm theo}h^{2}$.
DM direct detection Half Gaussian $60\,\mathrm{MeV}<m_{\rm{DM}}<1\,\mathrm{GeV}$ (DarkSide-50 Agnes et al. (2023)), $160\,\mathrm{MeV}<m_{\rm{DM}}<1\,\mathrm{GeV}$ (CRESST Abdelhameed et al. (2019); Angloher et al. (2023)).
CMB Half Gaussian $\langle\sigma v\rangle^{\text{Planck}}_{\text{CMB}}$ for 95% C.L. Aghanim et al. (2020a)
DM Self Interaction Half Gaussian $\sigma_{\rm{DM},\rm{DM}\to\rm{DM},\rm{DM}}/m_{\text{DM}}<1.0~{}{\rm cm^{2}/g}$ Randall et al. (2008)
$\triangle N_{\rm eff}$ Half Gaussian $\triangle N_{\rm eff}<0.17$ for 95% C.L. Aghanim et al. (2020a)
BBN Conditions $\tau_{\text{MED}}\leq 10^{5}$ s Krnjaic (2016)
Mediator production Conditions See Refs. Andreev et al. (2025, 2024); Adachi et al. (2024); Lees et al. (2016); Andreev et al. (2021) for details.
Table 1: Experimental constraints and likelihood distributions used in our analysis.
In this section, we discuss the constraints and their corresponding likelihoods in our study. Focusing on DM particles with masses between MeV and GeV, we explore the parameter space of these three benchmark models, see Eq. (8), (9), and (10), subject to the constraints summarized in Table 1. We use FeynRules Alloul et al. (2014) to implement the Lagrangian and CalcHEP Belyaev et al. (2013) to calculate the cross-sections.
@@
where $\mu_{\text{theo}}$ represents the theoretical prediction and $\mu_{\rm exp}$ denotes the experimental central value. The total uncertainty $\sigma_{\rm tot.}$ is given by $\sigma_{\rm tot.}=\sqrt{\sigma_{\rm theo}^{2}+\sigma_{\rm exp}^{2}}$. For a Half-Gaussian likelihood, where no signal is expected, we set $\mu_{\rm exp}=0$. In the following, we outline relevant constraints in Table 1 and their associated likelihoods.
@@
    We use the Boltzmann solver MicrOMEGAs Belanger et al. (2021) to compute the predicted DM relic density $\Omega^{\text{theo}}_{\text{DM}}h^{2}$. For likelihood, we use the PLANCK 2018 data Aghanim et al. (2020a) with the central value and statistical error from Table 1. In addition to statistical error, we include a systematic error of $\sigma_{\rm{sys}}=10\%$, accounting for uncertainties in the Boltzmann solver and the entropy table.
@@
    For sub-GeV DM particles scattering with nuclei, the low energy deposition ($v_{\rm DM}\approx 10^{-3}$) can only be detected in lower-threshold experiments, such as DarkSide-50 Agnes et al. (2023) and CRESST-III Abdelhameed et al. (2019); Angloher et al. (2023), which are most sensitive in the DM mass range $60\ \text{MeV}<m_{\rm DM}<1\ \text{GeV}$ for DM-nucleon interactions. Leptophilic interactions do not involve tree-level DM-quark elastic scattering, but DM can couple to nucleons via photon and lepton loops. We use the general lepton current ($\bar{l}\Gamma_{l}l$) to calculate the theoretical spin-independent DM-nucleon scattering cross-section $\sigma_{p}^{\text{SI}}$.
    For a scalar lepton current ($\Gamma_{l}=1$), the scalar current does not couple to a vector current, thus DM-quark interactions arise only at the two-loop level for both $\mathcal{L}_{SS}$ and $\mathcal{L}_{FS}$. We compute $\sigma_{p}^{\text{SI}}$ using Eq. (B1) and (B5) from Ref. Abdughani et al. (2022).
@@
    In addition to DM-nucleon scattering, sub-GeV DM can scatter off electrons, producing detectable single-electron ionization signals. However, recent results from XENON1T Aprile et al. (2019), SENSEI Barak et al. (2020), and DarkSide-50 Agnes et al. (2018) place stringent limits on such interactions. While electrophilic DM couples to electrons at tree level, muonphilic DM lacks this interaction and relies on suppressed loop-level contributions. As no DM-electron scattering signal has been observed, these experiments impose strong constraints on the parameter space of electrophilic DM.
    We compute the DM-free electron scattering cross-section $\bar{\sigma}_{e}$ using the CalcHEP Belyaev et al. (2013) package. Assuming a direct coupling between DM and electrons, we adopt the model-independent parameterization from Ref. Essig et al. (2012), with the reference cross-section defined as
@@
    The squared matrix element averaged over initial and summed over final spins with 3-momentum transfer $q$ is denoted as $\overline{\left|\mathcal{M}_{{\rm DM}e}(q)\right|^{2}}$. For models with a light mediator ($m_{\rm{MED}}\ll\rm{keV}$), the form factor is $F_{\rm{DM}}(q)=(\alpha m_{e}/q)^{2}$ Essig et al. (2012, 2020). In our case, with $m_{\rm{MED}}\gg\alpha m_{e}$, we take $F_{\rm{DM}}(q)=1$.
    To better illustrate the impact of DM-electron scattering constraints, we do not include them in the likelihood analysis. These limits mainly constrain electrophilic DM, as muonphilic DM interacts with electrons only via suppressed loop-level Feynman diagrams. Moreover, they are insensitive to the resonance region. Instead, we show these constraints directly in the $(m_{\rm DM},\bar{\sigma}_{e})$ plane, where their effect on the excluded parameter space is more clearly visualized. Detailed results and discussions are presented in Sec. IV.
  • CMB constraints
    Although DM can annihilate into either two leptons ($s$-channel) or four leptons ($t$-channel with two mediators’ decay), CMB constraints are sensitive only to energy injection. Therefore, we apply the same upper limits from Planck 2018 data Aghanim et al. (2020a) for both two-body and four-body lepton final states at the same DM mass. The DM annihilation cross-section during the CMB era is denoted as $\langle\sigma v\rangle_{\text{CMB}}$, and its velocity-average $\langle\sigma v\rangle_{\rm{CMB}}$ is
@@
    The annihilation cross section $\sigma v$ is computed using the CalcHEP Belyaev et al. (2013). The DM velocity at recombination epoch $v_{\rm CMB}$ is derived as
@@
    For CMB likelihood, we use a Half-Gaussian distribution
@@
    where $\langle\sigma v\rangle^{\text{Planck},95\%}_{\text{CMB}}$ is the upper limit on the cross section for a given DM mass at the $95\%$ confidence level, as reported by Planck 2018 data Aghanim et al. (2020a).
@@
  • $N_{\rm eff}$ constraint
    A light mediator may contribute to the relativistic degrees of freedom $N_{\rm eff}$. We incorporate this effect by implementing a half Gaussian likelihood with a $95\%$ confidence level upper limit $\Delta N_{\rm eff}<0.17$ Aghanim et al. (2020a). This constraint effectively rules out the parameter space where both $m_{\rm MED}$ and $m_{\rm DM}$ are below a few MeV.
  • BBN
    Mediator decay into SM particles may disrupt BBN history. However, in this work, we assume the mediator decays only into leptons. To avoid violating BBN constraints, we impose the conservative condition $\tau_{\text{MED}}\leq 10^{5}\,\text{s}$ Krnjaic (2016).
  • Mediator production
    For mediator particles in the MeV to GeV mass range, the most sensitive beam-dump experiments for detecting leptonphilic mediator production are NA64 Andreev et al. (2025, 2024), Belle II Adachi et al. (2024), and BABAR Lees et al. (2016). However, both theoretical models and experimental constraints are categorized based on electronphilic and muonphilic mediators. Since each experiment aims different mediator types and employs distinct analysis methods, we apply conditional cuts rather than likelihoods, ensuring consistency with their respective predictions and measurements.
[…]
    For $m_{\rm MED}>2m_{\mu}$, the mediator can decay into muons, and Belle II Adachi et al. (2024) and BABAR Lees et al. (2016) impose the strongest constraints. The mediator is produced via $e^{+}e^{-}\to{\rm MED}~{}\mu^{+}\mu^{-}$, and successively decays as ${\rm MED}\to 2\mu$. This process allows us to probe $g_{f}$ as a function of the mediator mass. For $m_{\rm MED}<1\,\mathrm{GeV}$, both BABAR and Belle II require $g_{f}\lesssim 10^{-3}$ at 95% confidence level.
@@ IV Allowed parameter space
Figure 2: The 95% confidence regions of three interactions projected in the $m_{\text{DM}}$ vs. $m_{\text{MED}}$ plane. The upper panels show the electrophilic DM scenario, and the lower panels show the muonophilic DM scenario. The red and orange regions correspond to the resonance scenario, while the gray region indicates forbidden annihilation mechanisms. The blue and green regions represent the parameter space where $m_{\rm{DM}}>m_{\rm{MED}}$.
@@
We use emcee Foreman-Mackey et al. (2013), a MCMC sampler, to explore the parameter space with the likelihood functions from Sec. III.2. We run 20 Markov chains for each model to ensure the coverage of the four-dimensional parameter space defined in Eq. (16). To present our results, we apply the Profile Likelihood method to eliminate nuisance parameters and project the results onto a two-dimensional plane. Assuming a nearly Gaussian likelihood, we define the two-dimensional $95\%$ confidence region as $\delta\chi^{2}<5.99$.
@@
Fig. 2 presents the $95\%$ confidence region in the $(m_{\rm DM},m_{\rm MED})$ plane, with electrophilic and muonphilic DM shown in the upper and lower panels, respectively. The red and orange regions indicate resonance annihilation ($m_{\rm MED}\approx 2m_{\rm DM}$), where DM satisfies thermal relic density and CMB limits, independent of $s$-, $p$-, or $(s+p)$-wave processes. However, CMB data exclude the resonance $s$-wave annihilation parameter space for $m_{\rm DM}<10\,\mathrm{MeV}$, as seen in the upper left panel. Additionally, in electrophilic scenarios, the lower bound on $m_{\rm{MED}}$ is set by constraints from $\Delta N_{\rm eff}$ measurements.
The gray region represents the forbidden DM annihilation area $m_{\rm MED}\approx m_{\rm DM}$, where DM can annihilate into mediator pairs in the early universe, yielding the correct relic density. However, at present-day Galactic temperatures, the annihilation of non-relativistic DM is kinetically forbidden, even in high-density environments like dwarf galaxies. As a result, DM in this region hardly produce detectable annihilation signals in indirect detection experiments.
@@
For $s$-wave annihilation, the cross-section is constant, resulting in stringent CMB constraints and leaving only the resonance and forbidden regions viable. In contrast, $p$-wave annihilation escapes CMB constraints due to velocity suppression, allowing a broader parameter space. The parameter space with $m_{\rm{DM}}>m_{\rm{MED}}$ becomes particularly interesting for the $(s+p)$-wave mechanism. We can parameterize the velocity dependence of the annihilation cross section as $\langle\sigma v\rangle\simeq a+bv^{2}$, assuming lower velocities. Ref. Abdughani et al. (2022) shows that in a vector mediator scenario characterized by the Lagrangian $\mathcal{L}_{FV}$, the cross section $\langle\sigma v\rangle_{2f}$ for the process ${\rm DM}\,{\rm DM}\to f\bar{f}$ yields a ratio $a/b\approx\mathcal{O}(1)$. Given the extremely low DM velocities during the CMB epoch ($v\sim 10^{-6}$ to $10^{-8}$), the cross section $\langle\sigma v\rangle_{2f}$ is dominated by the $s$-wave component, which excludes the parameter space $m_{\rm{DM}}\gtrsim m_{\rm{MED}}$ based on CMB data. In contrast, for the process ${\rm DM}\,{\rm DM}\to{\rm MED}\,{\rm MED}\to 4f$, the ratio $a/b\ll 1$ indicates that the cross section $\langle\sigma v\rangle_{4f}$ behaves like a $p$-wave.
@@
Figure 3: The $95\%$ allowed region for the DM-electron scattering cross-section $\bar{\sigma}_{e}$ as a function of $m_{\text{DM}}$. The region labeled as non-resonance region includes the forbidden and $m_{\rm DM}>m_{\rm MED}$ scenarios, as defined in Fig. 2. The form factor is taken as heavy mediator limit, $F_{\rm DM}(q)=1$. The region above the black solid line is excluded by the null results from the XENON1T Aprile et al. (2019) DM search experiment.
@@
Considering direct detection for DM-electron interactions, the electrophilic and muonphilic scenarios yield very different predictions. In the electrophilic case, DM couples to electrons through a tree-level diagram, which significantly enhances the DM-free electron scattering cross-section $\bar{\sigma}_{e}$, making it accessible for current direct detection experiments. In contrast, in the muonphilic case, the absence of a tree-level contribution to $\bar{\sigma}_{e}$ leads to substantial suppression from loop-level processes.

### v2

@@ III.1 Benchmark models
In Ref. Abdughani:2021oit, all representative interactions for muonphilic DM models were systematically classified. These models include DM and mediator (MED) particles, which can be scalar, fermion, or vector. To prevent DM decay, we introduced a $Z_{2}$ symmetry and constructed 16 renormalizable effective models with $Z_{2}$-even mediators, as summarized in Table II of Ref. Abdughani:2021oit.
In this study, we first perform a preliminary scan of these 16 models and categorize them into three groups based on their annihilation characteristics: $s$-wave, $p$-wave, and ss+$p$-wave. Moreover, we find that models exhibiting the same annihilation behavior share highly similar surviving parameter spaces after incorporating the experimental constraints listed in Table 1. Given these findings, we select three benchmark models as the focus of our analysis and presentation: scalar DM with a scalar mediator, Dirac DM with a scalar mediator, and Dirac DM with a vector mediator. The corresponding Lagrangians for these models are denoted as $\mathcal{L}_{SS}$, $\mathcal{L}_{FS}$, and $\mathcal{L}_{FV}$, respectively. The detailed discussions of these models is presented in the following section.
For DM particles promptly annihilating into detectable $\gamma$-rays with energy $\mathcal{O}(10\ \text{MeV})$, their mass must be in the MeV to GeV range. Additionally, studies Matsumoto:2018acr; Chen:2024njd; Chen:2018vkr indicate that the constraints of the thermal relic density requires sub-GeV DM annihilation via a mediator with mass below GeV. For leptophilic mediators, such a light particle can only decay into muons or electrons. Therefore, to focus on leptophilic mediators, we individually study two annihilation channels: (i) ${\rm DM}+{\rm DM}\to e^{+}+e^{-}$ or ${\rm DM}+{\rm DM}\to 4e$, and (ii) ${\rm DM}+{\rm DM}\to\mu^{+}+\mu^{-}$ or ${\rm DM}+{\rm DM}\to 4\mu$. Based on these motivations, we list three representative interactions.
@@
    where $f$ can be either $e^{\pm}$ or $\mu^{\pm}$ in this work. The dimensional coupling between $S$ and $\phi$ is $M_{D}$, while $g_{f}$ denotes the coupling between the mediator and the fermions. This interaction suggests that DM annihilation proceeds via $s$-wave, with a cross section around $\mathcal{O}(10^{-26})$ cm-3s-1 to produce the correct relic density $\Omega h^{2}\approx 0.1$. However, such $s$-wave annihilation with DM masses below $1\,\mathrm{GeV}$ may inject excessive energy, potentially distorting the temperature and polarization spectra of the CMB Matsumoto:2018acr; Chen:2024njd; Wang:2025tdx. Therefore, CMB observations impose stringent constraints on the coupling constants $g_{D}$ and $g_{f}$, particularly for sub-GeV DM $s$-wave annihilation.
@@ III.2 Experimental Constraints
Observable Likelihood type Experimental data
Relic abundance Gaussian $\Omega_{\rm{DM}}^{\rm exp}h^{2}=0.1193\pm 0.0014$ Planck:2018vyg; $\sigma_{\rm sys}=10\%\times\Omega_{\rm{DM}}^{\rm theo}h^{2}$.
DM direct detection Half-Gaussian $60\,\mathrm{MeV}<m_{\rm{DM}}<1\,\mathrm{GeV}$ (DarkSide-50 DarkSide-50:2023fcw), $160\,\mathrm{MeV}<m_{\rm{DM}}<1\,\mathrm{GeV}$ (CRESST CRESST:2019jnq; CRESST:2022lqw).
CMB Half-Gaussian $\langle\sigma v\rangle^{\text{Planck}}_{\text{CMB}}$ for 95% C.L. Planck:2018vyg
DM Self Interaction Half-Gaussian $\sigma_{\rm{DM},\rm{DM}\to\rm{DM},\rm{DM}}/m_{\text{DM}}<1.0~{\rm cm^{2}/g}$ Randall:2008ppe
$\triangle N_{\rm eff}$ Half-Gaussian $\triangle N_{\rm eff}<0.17$ for 95% C.L. Planck:2018vyg
BBN Conditions $\tau_{\text{MED}}\leq 10^{5}$ s Krnjaic:2015mbs
Mediator production Conditions See Refs. 2502.04053; 2409.10128; 2403.02841; 1606.03501; NA64:2021xzo for details.
Table 1: Summary of observables, likelihood function types, and experimental data used in this work.
In this section, we summarize the experimental and observational data, and explain how they are incorporated into the likelihood functions used in our MCMC analysis. Focusing on DM particles with masses between MeV and GeV, we explore the parameter space of these three benchmark models, see Eq. (8), (9), and (10), subject to the constraints summarized in Table 1. We use FeynRules Alloul:2013bka to implement the Lagrangian and CalcHEP Belyaev:2012qa to calculate the cross sections.
@@
where $\mu_{\text{theo}}$ represents the theoretical prediction and $\mu_{\rm exp}$ denotes the experimental central value. The total uncertainty $\sigma_{\rm tot.}$ is given by $\sigma_{\rm tot.}=\sqrt{\sigma_{\rm theo}^{2}+\sigma_{\rm exp}^{2}}$. For a Half-Gaussian likelihood, where no signal is expected, we set $\mu_{\rm exp}=0$. In the following, we outline relevant constraints in Table 1 and their associated likelihood functions.
@@
    We use the Boltzmann solver MicrOMEGAs Belanger:2020gnr to compute the predicted DM relic density $\Omega_{\rm DM}^{\rm theo}h^{2}$. We implement a Gaussian likelihood function adopting the central value and statistical error from the Planck 2018 measurement Planck:2018vyg. In addition, we include a systematic uncertainty of $\sigma_{\rm sys}=10\%$ in the relic density calculation. This accounts for theoretical uncertainties associated with the effective number of relativistic degrees of freedom, $g_{\star}(T)$ and $g_{*S}(T)$, entering the Boltzmann equation during freeze-out. In the sub-GeV regime, freeze-out can occur near the QCD phase transition, where non-perturbative effects lead to $\mathcal{O}(10\%)$ variations in $g_{\star}$ and $g_{*S}$ among different evaluations Drees:2015exa. Such differences translate directly into comparable shifts in the predicted relic density. To remain conservative, we assume the maximal $\sim 10\%$ uncertainty reported in the literature Fowlie:2013oua; Kowalska:2014hza; Saikawa:2020swg; Watanabe:2025pvc.
@@
[…]
    For a scalar lepton current ($\Gamma_{l}=1$), the scalar current does not couple to a vector current, thus DM-quark interactions arise only at the two-loop level for both $\mathcal{L}_{SS}$ and $\mathcal{L}_{FS}$. We compute $\sigma_{p}^{\text{SI}}$ using Eq. (B1) and (B5) from Ref. Abdughani:2021oit.
@@
    In addition to DM-nucleon scattering, sub-GeV DM can scatter off electrons, producing detectable single-electron ionization signals. However, recent results from XENON1T XENON:2019gfn, DarkSide-50 DarkSide:2022knj, DAMIC-M DAMIC-M:2023gxo, and SENSEI SENSEI:2020dpa place stringent limits on such interactions. While electrophilic DM couples to electrons at tree level, muonphilic DM lacks this interaction and relies on suppressed loop-level contributions. As no DM-electron scattering signal has been observed, these experiments impose strong constraints on the parameter space of electrophilic DM.
    We compute the DM-free electron scattering cross section $\bar{\sigma}_{e}$ using the CalcHEP Belyaev:2012qa package. Assuming a direct coupling between DM and electrons, we adopt the model-independent parameterization from Ref. Essig:2011nj, with the reference cross section defined as
@@
    The squared matrix element averaged over initial and summed over final spins with 3-momentum transfer $q$ is denoted as $\overline{\left|\mathcal{M}_{{\rm DM}e}(q)\right|^{2}}$. For models with a light mediator ($m_{\rm{MED}}\ll\rm{keV}$), the form factor is $F_{\rm{DM}}(q)=(\alpha m_{e}/q)^{2}$ Essig:2011nj; Essig:2019xkx. In our case, with $m_{\rm{MED}}\gg\alpha m_{e}$, we take $F_{\rm{DM}}(q)=1$.
    To better illustrate the impact of DM-electron scattering constraints, we do not include these limits in the likelihood function, since they primarily affect electrophilic DM. Instead, we display them directly in the $(m_{\rm DM},\bar{\sigma}_{e})$ plane, where their effect on the excluded parameter space is more clearly visualized. Detailed results and discussions are presented in Sec. IV.
  • CMB
    DM can annihilate either directly into two leptons ($s$-channel) or into four leptons through mediator decays ($t$-channel). CMB anisotropies are sensitive only to the total electromagnetic energy deposited in the thermal plasma, which can be parameterized by the efficiency factor $f_{\rm eff}$ multiplying $\langle\sigma v\rangle/m_{\rm DM}$. While in principle $f_{\rm eff}$ differs somewhat between two-body and four-body leptonic final states, the variations are $\mathcal{O}(1)$ Slatyer:2015jla and do not qualitatively affect the likelihood function in the sub-GeV mass range relevant to this work. Therefore, we simply apply the updated Planck 2018 limits Planck:2018vyg to both channels.
    The DM annihilation cross section during the CMB era is denoted as $\langle\sigma v\rangle_{\text{CMB}}$, and its velocity-average $\langle\sigma v\rangle_{\rm{CMB}}$ is
@@
    The annihilation cross section $\sigma v$ is computed using the CalcHEP Belyaev:2012qa. The DM velocity at recombination epoch $v_{\rm CMB}$ is derived as
@@
    For CMB likelihood function, we use a Half-Gaussian distribution with Chi-squared defined as
@@
    where $\langle\sigma v\rangle^{\text{Planck},95\%}_{\text{CMB}}$ is the upper limit on the cross section for a given DM mass at the $95\%$ confidence level, as reported by Planck 2018 data Planck:2018vyg.
@@
    Since $g_{f}$ can be suppressed by DM direct detection or light mediator searches, $g_{D}$ shall be sizable to ensure the correct relic abundance. As a result, with a large $g_{D}$, the DM self-interaction cross section $\sigma_{\text{DM,DM}\to\text{DM,DM}}$ may become significant, potentially exceeding the upper limit set by the Bullet Cluster, $\sigma_{\text{DM,DM}\to\text{DM,DM}}/m_{\text{DM}}<1.0\,\text{cm}^{2}/\text{g}$ Randall:2008ppe, where the DM incident velocity is $4000\,\text{km/s}$.
  • $N_{\rm eff}$
    A light mediator may contribute to the relativistic degrees of freedom $N_{\rm eff}$. We incorporate this effect by implementing a half Gaussian likelihood function with a $95\%$ confidence level upper limit $\Delta N_{\rm eff}<0.17$ Planck:2018vyg. This constraint effectively rules out the parameter space where both $m_{\rm MED}$ and $m_{\rm DM}$ are below a few MeV.
  • BBN
    Mediator decay into SM particles may disrupt BBN thermal history. In this work, we assume the mediator decays only into leptons. To secure a successful BBN thermal history, we impose the conservative condition $\tau_{\text{MED}}\leq 10^{5}\,\text{s}$ Krnjaic:2015mbs.
  • Mediator production
    For mediator particles in the MeV to GeV mass range, the most sensitive beam-dump experiments for detecting leptonphilic mediator production are NA64 2502.04053; 2409.10128, Belle II 2403.02841, and BABAR 1606.03501. However, both theoretical models and experimental constraints are categorized based on electronphilic and muonphilic mediators. Since each experiment aims at different mediator types and employs distinct analysis methods, we apply conditional cuts to ensure consistency with their respective predictions and measurements.
@@
    For $m_{\rm MED}>2m_{\mu}$, the mediator can decay into muons, and Belle II 2403.02841 and BABAR 1606.03501 impose the strongest constraints. The mediator is produced via $e^{+}e^{-}\to{\rm MED}~\mu^{+}\mu^{-}$, and successively decays as ${\rm MED}\to 2\mu$. This process allows us to probe $g_{f}$ as a function of the mediator mass. For $m_{\rm MED}<1\,\mathrm{GeV}$, both BABAR and Belle II require $g_{f}\lesssim 10^{-3}$ at 95% confidence level.
@@ IV Allowed parameter space
Figure 2: The $2\sigma$ allowed regions ($\delta\chi^{2}<5.99$ for two degrees of freedom) in the $(m_{\text{DM}},m_{\text{MED}})$ plane for the three benchmark interactions. The upper panels correspond to the electrophilic DM scenario, while the lower panels correspond to the muonophilic DM scenario. Colored dots represent surviving parameter points from the MCMC scan. Red/orange indicate resonance regions, gray corresponds to forbidden annihilation, and blue/cyan mark the parameter space where $m_{\rm DM}>m_{\rm MED}$.
@@
We use emcee Foreman-Mackey:2012any, a MCMC sampler, to explore the parameter space with the likelihood functions from Sec. III.2. We run 20 Markov chains for each model to ensure the coverage of the four-dimensional parameter space defined in Eq. (16). To present our results, we apply the Profile Likelihood method to eliminate nuisance parameters and project the results onto a two-dimensional plane. Assuming a nearly Gaussian likelihood, we adopt the $2\sigma$ (95% C.L.) allowed region as $\delta\chi^{2}<5.99$ for two degrees of freedom.
@@
Fig. 2 shows the $2\sigma$ allowed region in the $(m_{\rm DM},m_{\rm MED})$ plane, with electrophilic and muonophilic DM depicted in the upper and lower panels, respectively. The colored dots represent parameter points satisfying total Chi-squared $\delta\chi^{2}<5.99$. Colors distinguish the surviving points based on their dominant annihilation mechanisms: red/orange for resonance, gray for forbidden regions, and blue/cyan for $m_{\rm DM}>m_{\rm MED}$. Specifically, red and blue denote the electrophilic DM (upper panels), while orange and cyan correspond to muonophilic DM (lower panels). The red and orange regions indicate resonance annihilation ($m_{\rm MED}\approx 2m_{\rm DM}$), where DM satisfies thermal relic density and CMB limits, independent of $s$-, $p$-, or $(s+p)$-wave processes. However, CMB data exclude the resonance $s$-wave annihilation parameter space for $m_{\rm DM}<10\,\mathrm{MeV}$, as seen in the upper left panel. Additionally, in electrophilic scenarios, the lower bound on $m_{\rm{MED}}$ and $m_{\rm{DM}}$ are set by constraints from $\Delta N_{\rm eff}$ measurements.
The gray region corresponds to the kinematic threshold regime where $m_{\rm MED}\approx m_{\rm DM}$. In the early universe, when DM was semi-relativistic, annihilation into mediator pairs was possible and could efficiently deplete the abundance to yield the observed relic density. In contrast, in the present-day Galactic environment where DM is non-relativistic ($v\sim 10^{-3}$), the available kinetic energy is far too small to overcome the mass threshold. As a result, annihilation into mediator pairs is forbidden today, and this parameter region does not lead to observable indirect detection signals, even in dense DM systems such as dwarf spheroidal galaxies.
@@
For $s$-wave annihilation, the cross section is constant, resulting in stringent CMB constraints and leaving only the resonance and forbidden regions viable. In contrast, $p$-wave annihilation escapes CMB constraints due to velocity suppression, allowing a broader parameter space. The parameter space with $m_{\rm{DM}}>m_{\rm{MED}}$ becomes particularly interesting for the $(s+p)$-wave mechanism. We can parameterize the velocity dependence of the annihilation cross section as $\langle\sigma v\rangle\simeq a+bv^{2}$, assuming lower velocities. Ref. Abdughani:2021oit shows that in a vector mediator scenario characterized by the Lagrangian $\mathcal{L}_{FV}$, the cross section $\langle\sigma v\rangle_{2f}$ for the process ${\rm DM}\,{\rm DM}\to f\bar{f}$ yields a ratio $a/b\approx\mathcal{O}(1)$. Given the extremely low DM velocities during the CMB epoch ($v\sim 10^{-6}$ to $10^{-8}$), the cross section $\langle\sigma v\rangle_{2f}$ is dominated by the $s$-wave component, which excludes the parameter space $m_{\rm{DM}}\gtrsim m_{\rm{MED}}$ based on CMB data. In contrast, for the process ${\rm DM}\,{\rm DM}\to{\rm MED}\,{\rm MED}\to 4f$, the ratio $a/b\ll 1$ indicates that the cross section $\langle\sigma v\rangle_{4f}$ behaves like a $p$-wave.
@@
Figure 3: The $2\sigma$ allowed region for the DM-electron scattering cross section $\bar{\sigma}_{e}$ as a function of $m_{\text{DM}}$. The region labeled as non-resonance region includes the forbidden and $m_{\rm DM}>m_{\rm MED}$ scenarios, as defined in Fig. 2. The form factor is taken as heavy mediator limit, $F_{\rm DM}(q)=1$. The region above the solid lines is excluded by the null results from the XENON1T XENON:2019gfn, DarkSide-50 DarkSide:2022knj, and DAMIC-M DAMIC-M:2023gxo DM search experiments.
@@
Considering direct detection for DM-electron interactions, the electrophilic and muonphilic scenarios yield very different predictions. In the electrophilic case, DM couples to electrons through a tree-level diagram, which significantly enhances the DM-free electron scattering cross section $\bar{\sigma}_{e}$, making it accessible for current direct detection experiments. In contrast, in the muonphilic case, the absence of a tree-level contribution to $\bar{\sigma}_{e}$ leads to substantial suppression from loop-level processes.